OpenStudy (anonymous):

The Fun Guys game rental store charges an annual fee of $5 plus $5.50 per game rented. The Game Bank charges an annual fee of $20 plus $2.50 per game. For how many game rentals will the cost be the same at both stores? Games= What is that cost? dollars=

OpenStudy (anonymous):

Cost at Fun Guys = Cost at Game Bank 5 + 5.5x = 20 + 2.5x x is the number of game rentals. Solve for x

OpenStudy (anonymous):

2x=15

OpenStudy (anonymous):

3x = 15

OpenStudy (anonymous):

5.5 - 2.5 = 3

OpenStudy (anonymous):

sry bb im bad at math

OpenStudy (anonymous):

looks like an easy finish!

OpenStudy (anonymous):

whats da cost thou?

OpenStudy (anonymous):

well, how many games is it?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

tell me this, and then take that number and put it into one of the equations

OpenStudy (anonymous):

5 + 5.5 * x = 20 + 2.5 * x 5.5 * x - 2.5 * x = 20 - 5 3 * x = 15 x = 5 Games

OpenStudy (anonymous):

oh see im bad at math

OpenStudy (anonymous):

ooh. no no! you were at 3x=15 x is the number of games to make the prices equal!

OpenStudy (anonymous):

3x means 3 times x =15

OpenStudy (anonymous):

5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so, if both stores the price will be the same if people rent 5 games, you take that 5 and plug it into either equation you want, because they are equal!

OpenStudy (anonymous):

ok thank ya ok

OpenStudy (anonymous):

and pay 32.5

OpenStudy (anonymous):

see the original question. FUN GUYS is 5 dollars +5.50 per game. you are looking at the cost for 5 games so you need 5 times 5.50 because 5.50 is for only 1 game. Answer = 5 + 5*5.50

OpenStudy (anonymous):

hazem_wardet thx 4 the help bru and jeremyggg its cool to help another GUY out hehehehe

OpenStudy (anonymous):

itz a pic of my gf bruhs

OpenStudy (anonymous):

try and set up the same thing for the other equation if you want the practice.

OpenStudy (anonymous):

haha i dont really look at the pictures anyway so i wasn't assuming!

OpenStudy (anonymous):

nah im good,hazem was he pm me ;) lol

OpenStudy (anonymous):

ok! good luck with your math!

OpenStudy (anonymous):

lol thnx bro